# CI/CD Configuration Practices in Open-Source Android Apps: An Empirical Study

### v1

@@ 8. Conclusion
This paper provides a comprehensive analysis of CI/CD practices in Android apps, revealing the complexities and challenges developers face due to a lack of standardization in CI/CD configurations across the available services. Our empirical results suggest a significant focus on the setup phase in most CI/CD pipelines, with limited emphasis on testing and deployment. The findings also indicate that active maintenance, though infrequent, correlates with factors like project maturity and community engagement. Overall, our study highlights the need for standardized practices to improve interoperability and suggests that automation and AI-driven solutions could significantly enhance CI/CD processes. Furthermore, it is recommended to establish adaptable frameworks and open-source tools to overcome resource constraints, particularly in testing and deployment. Addressing these aspects could facilitate smoother adoption and maintenance of CI/CD pipelines in mobile development, ultimately leading to more efficient and effective software delivery.
In the future, we aim to seek feedback from software developers on our empirical findings to guide future research into the areas of highest demand for automation in CI/CD processes. We also plan to study how deficiencies in CI/CD pipelines are linked to mobile app success, particularly user perception like app ratings in app stores.

### v2

@@ 8. Conclusion
This paper provides a comprehensive analysis of CI/CD practices in Android apps, revealing the complexities and challenges developers face due to a lack of standardization in CI/CD configurations across the available services. Our empirical results suggest a significant focus on the setup phase in most CI/CD pipelines, with limited emphasis on testing and deployment. The findings also indicate that active maintenance, though infrequent, correlates with factors like project maturity and community engagement. Our findings agree with prior research on general-purpose CI/CD practices regarding configuration complexity, but are somewhat inconsistent with certain aspects of configuration change frequency and themes. Overall, our study highlights the need for standardized practices to improve interoperability and suggests that automation and AI-driven solutions could significantly enhance CI/CD processes. Furthermore, it is recommended to establish adaptable frameworks and open-source tools to overcome resource constraints, particularly in testing and deployment. Addressing these aspects could facilitate smoother adoption and maintenance of CI/CD pipelines in mobile development, ultimately leading to more efficient and effective software delivery.
In the future, we aim to seek feedback from software developers on our empirical findings to guide future research into the areas of highest demand for automation in CI/CD processes. We also aim to expand our analyses in future work to include a broader range of application domains to support more generalizable conclusions and to explore domain-specific patterns in CI/CD configuration and maintenance practices. Moreover, we plan to study how deficiencies in CI/CD pipelines are linked to mobile app success, particularly user perception (e.g., app ratings in app stores).
@@ Acknowledgment
9. Data Availability
The replication package of our study (including the data, code, and raw results) is available on Figshare (Ghaleb et al., 2024).